OpenStudy (anonymous):

You are making poster board boundary markers in the shape of square pyramids for a relay race. If you are including a 10-inch square as the bottom and you want the slant height to be 12inches, how much poster board do you need?

OpenStudy (anonymous):

340in.2 360in.2 400in.2 384in.2

OpenStudy (anonymous):

@karatechopper , @annas , @seashell anyone help plzz

OpenStudy (anonymous):

There are 4 triangular sides (one for each side of the square), so we multiply the area of ONE triangle by 4 112 in2 × 4 = 448 in2 Total surface area is 196 in2 + 448 in2 = 644 in2 --->644 in2

OpenStudy (anonymous):

it cant be 644

OpenStudy (anonymous):

the 2 i put after in. is small like this

OpenStudy (anonymous):

@seashell plz answer me

OpenStudy (anonymous):

are you still here

OpenStudy (anonymous):

sorry i wasnt on

OpenStudy (anonymous):

can you help me now

OpenStudy (anonymous):

yeah :) hold on... im rechecking :)

OpenStudy (anonymous):

you have 1 square, area 100. Then each triangle has area (0.5 x 10 x 12)=60, and there are four. so 100 + 4(60)= 100+240 = 340

OpenStudy (anonymous):